OpenStudy (anonymous):

medal included!! Solve. 18 = – 2m + 6 – m m =

OpenStudy (anonymous):

-4

geerky42 (geerky42):

Would be better if you explain how you got that answer.

OpenStudy (anonymous):

is that right people

OpenStudy (anonymous):

18=-2m+6-m Combine like terms 18=-3m+6 Subtract 6 from both sides 12=-3m Divide both sides by -3 -4=m

OpenStudy (anonymous):

thank you